OpenStudy (anonymous):

Earl has 28 coins, all dimes and quarters. How many of each does he have, if their total value is $5.20?

OpenStudy (anonymous):

d+q=5.20 .10d+.25q=5.20 sub method

OpenStudy (anonymous):

wait! d+q=28 sorry

OpenStudy (anonymous):

now you sub. lets do dimes. d=28-q so sub it .10(28-q)+.25q=5.20 2.80-.10q+.25q=5.20 .15q=2.40 q=16 soooo then dimes is 12

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yes, thank u